# A Survey on Self-Evolution of Large Language Models

### v1

@@ 4.2.2 Negative
In addition to acquiring positive solutions, recent research illustrates that LLMs can benefit from negative ones for self-improvement Yang et al. (2023b). This strategy is analogous to trial and error in human behavior when learning skills. This section summarises typical methods of gaining negative solutions to assist in self-evolution.
@@
Self-Reward, SPIN Yuan et al. (2024); Chen et al. (2024) updates the model by comparing the answers of high and low scores. Similarly, GRATH Chen and Li (2024) generates both correct and incorrect answers. It then trains the model by comparing these two answers. Self-Contrast Zhang et al. (2024c) contrasts the differences and summarizes these discrepancies into a checklist that could be used to re-examine and eliminate discrepancies. In ETO Song et al. (2024), the model interacts with the embodied environment to complete tasks and optimizes from the failure solutions. A3Yang et al. (2024c) improves ETO by adding rationale after each action for solving tasks. STE Wang et al. (2024b) implements trial and error where the model solves the tasks with unfamiliar tools. It learns by analyzing failed attempts to improve problem-solving strategies in future tasks. More recently, COTERRORSET Tong et al. (2024) obtains incorrect solutions generated by PALM-2 and proposes mistake tuning, which requires the model to void making mistakes.
@@
Some methods add perturbation to generate harmful solutions Yang et al. (2023a); Liu et al. (2024a). Given a task, RLCD Yang et al. (2023a) curates both positive and negative instructions and generates positive and negative solutions. DLMA Liu et al. (2024a) gathers both positive and negative instructional prompts and subsequently produces corresponding positive and negative solutions.
Rather than harmful perturbation, incorporating negative context is another way. Ditto Lu et al. (2024a) adds negative persona characters to generate incorrect conversations. The model then learns from the negative conversations to evolve persona dialogue ability.
@@ 4.3 Feedback
${\mathcal{F}}^{t}=f^{{\mathcal{F}}}({\mathcal{T}}^{t},{\mathcal{Y}}^{t},{\mathcal{E}}^{t},\mathrm{M}^{t},\mathrm{ENV}),$ (4)
@@ 4.3.1 Model
Current studies demonstrate that LLMs can play well as a critic Zheng et al. (2024a). In the cycle of self-evolution, the model judges itself to acquire the feedback of the solutions.
One type of feedback is a score that indicates correctness. Self-Reward Yuan et al. (2024), LSX Stammer et al. (2023), and DLMA Liu et al. (2024a) rate their own solutions and output the scores via LLM-as-a-Judge prompting. Similar to that, SIRLC Pang et al. (2023) utilizes self-evaluation results of LLM as the reward for further reinforcement learning. Self-Alignment Zhang et al. (2024e) leverages the self-evaluation capability of an LLM to generate confidence scores on the factual accuracy of its outputs.
Another type provides a textual description, offering multi-dimensional information. To alter the distribution of the responses via supervised learning, CAI Bai et al. (2022) asks the model to critique its response according to a principle in the constitution. In contrast to supervised learning and reinforcement learning approaches, Self-Refine Madaan et al. (2023) allows the model to generate natural language feedback on its own output in a few-shot manner.
@@ 4.3.2 Environment
Another form of feedback comes from the environment, common in tasks where solutions can be directly evaluated. This feedback is precise and elaborate and can provide sufficient information for model updating. They may be derived from code interpreter Jiang et al. (2023); Chen et al. (2023c); Shinn et al. (2024), tool execution Qiao et al. (2024); Gou et al. (2023), the embodied environment Bousmalis et al. (2023); Xu et al. (2024b); Zhou et al. (2023b), and other LLMs or agents Wang et al. (2024e); Taubenfeld et al. (2024); Ulmer et al. (2024).
For code generation, Self-Debugging Chen et al. (2023c) utilizes execution results on test cases as part of feedback while SelfEvolve Jiang et al. (2023) receives the error message from the interpreter. Similarly, Reflexion Shinn et al. (2023) also obtains the run-time feedback from the code interpreter. It then further reflects to generate thoughts. This run-time feedback contains the trace-back information that can point out the key information for improved code generation.
Recently, methods endow tool-using ability to LLMs and agents. Executing tools leading to feedback in return Gou et al. (2023); Qiao et al. (2024); Song et al. (2024); Yang et al. (2024c); Wang et al. (2024b).
RoboCat Bousmalis et al. (2023) and SinViG Xu et al. (2024b) act in the robotic embodied environment. This type of feedback is precise and strong to guide self-evolution.
Communication feedback is common and effective in LLM-based multi-agent systems. Agents can correct and support each other, enabling co-evluation Wang et al. (2024e); Taubenfeld et al. (2024); Ulmer et al. (2024).
Figure 7: Experience refinement.
@@ 5.1.1 Metric-Based
By relying on feedback and pre-defined criteria, metric-based filtering improves the quality of the outputs Singh et al. (2023); Qiao et al. (2024); Ulmer et al. (2024); Wang et al. (2023b), ensuring the progressive enhancement of LLM capabilities through each iteration of refinement.
For example, ReSTEM Singh et al. (2023) incorporates a reward function to filter the dataset sampled from the current policy. The function provides binary rewards based on the correctness of the generated samples rather than a learned reward model trained on human preferences in ReST Gulcehre et al. (2023). AutoAct Qiao et al. (2024) leverages F1-score and accuracy as rewards for synthetic trajectories and collects trajectories with exactly correct answers for further training. Self-Talk Ulmer et al. (2024) measures the number of completed subgoals to filter the generated dialogues, ensuring that only high-quality data is used for training. To encourage diversity of the source instructions, Self-Instruct Wang et al. (2023b) automatically filters low-quality or repeated instructions using ROUGE-L similarity and heuristics before adding them to the task pool.
@@ 5.1.2 Metric-Free
Some methods seek filtering strategies beyond external metrics, making the process more flexible and adaptable. Metric-Free filtering typically involves sampling outputs and evaluating them based on internal consistency measures or other model-inherent criteria Huang et al. (2022); Weng et al. (2023); Chen et al. (2022). The filtering in Self-Consistency Wang et al. (2022) is based on the consistency of the final answer across multiple generated reasoning paths, with higher agreement indicating higher reliability. LMSI Huang et al. (2022) utilizes CoT prompting plus self-consistency for generating high-confidence self-training data.
Designing internal consistency measures that accurately reflect output quality can be challenging. Self-Verification Weng et al. (2023) allows the model to select the candidate answer with the highest interpretable verification score, calculated by assessing the consistency between the predicted and original condition values. For the code generation task, CodeT Chen et al. (2022) considers both the consistency of the outputs against the generated test cases and the agreement of the outputs with other code samples.
@@ 5.2 Correcting
Recent advancements in self-evolution have highlighted the significance of iterative self-correction, which enables models to refine their experiences. This section divides the methods employed into two categories: Critique-Based and Critique-Free correction. Critiques often serve as strong hints that include the rationale behind perceived errors or suboptimal outputs, guiding the model towards improved iterations.
@@ 5.2.1 Critique-Based
These methods rely on additional judging processes to draw the critiques of the experiences. Then, the experiences are refined based on the critiques. By leveraging either self-generated Madaan et al. (2023); Bai et al. (2022); Shinn et al. (2023); Lu et al. (2023) or environment-interaction generated critiques Gou et al. (2023); Jiang et al. (2023); Zhou et al. (2023b), the model benefits from detailed feedback for nuanced correction.
LLMs have demonstrated their ability to identify errors in their outputs. Self-Refine Madaan et al. (2023) introduces an iterative process in which the model refines its initial outputs conditioned on actionable self-feedback without additional training. To evolve from the correction, CAI Bai et al. (2022) generates critiques and revisions of its outputs in the supervised learning phase, significantly improving the initial model. Applied to an agent automating computer tasks, RCI Kim et al. (2023) improves its previous outputs based on the critique finding errors in the outputs.
Since weaker models may struggle significantly with the self-critique process, several approaches enable models to correct the outputs using critiques provided by external tools. CRITIC Gou et al. (2023) allows LLMs to revise the output based on the critiques obtained during interaction with tools in general domains. SelfEvolve Jiang et al. (2023) prompts an LLM to refine the answer code based on the error information thrown by the interpreter. ISR-LLM Zhou et al. (2023b) helps the LLM planner find a revised action plan by using a validator in an iterative self-refinement process.
@@ 5.2.2 Critique-Free
Contrary to critique-based, critique-free methods correct the experiences directly leveraging objective information Zelikman et al. (2022); Chen et al. (2023c, b); Gero et al. (2023). These methods offer the advantage of independence from nuanced feedback that critiques provide, allowing for corrections that adhere strictly to factual accuracy or specific guidelines without the potential bias introduced by critiques.
One group of critique-free methods modifies the experiences on the signal of whether the task was correctly resolved. Self-Taught Reasoner (STaR) Zelikman et al. (2022) proposes a technique that iteratively generates rationales to answer questions. If the answers are incorrect, the model is prompted again with the correct answer to generate a more informed rationale. Self-Debug Chen et al. (2023c) enables the model to perform debugging steps by investigating execution results from unit tests and explaining the code on its own.
Different from depending on the task-solving signal, other information produced during the solving process can be leveraged. IterRefinement Chen et al. (2023b) relies on a series of refined prompts that encourage the model to reconsider and improve upon its previous outputs without any direct critique. For information extraction tasks, Clinical SV Gero et al. (2023) grounds each element in evidence from the input and prunes inaccurate elements using supplied evidence.
@@ 6.1 In-Weight
Classical training paradigms in updating LLMs in weight encompass continuous pretraining Brown et al. (2020); Roziere et al. (2023), supervised fine-tuning Longpre et al. (2023), and preference alignment Ouyang et al. (2022); Touvron et al. (2023a). However, in the iterative training process of self-evolving, the core challenge lies in achieving overall improvement and preventing catastrophic forgetting, which entails refining or acquiring new capabilities while preserving original skills. Solutions to this challenge can be categorized into three main strategies: replay-based, regularization-based, and merging-based methods.
@@ 6.1.1 Replay-based
Replay-based methods reintroduce previous data to retain old knowledge. One is experience replay, which mixes the original and new training data to update LLMs Roziere et al. (2023); Yang et al. (2024c); Zheng et al. (2023); Lee et al. (2024); Wang et al. (2023a). For example, Rejection sampling Fine-Tuning (RFT) Yuan et al. (2023) and Reinforced Self-Training (ReST) Gulcehre et al. (2023); Aksitov et al. (2023) method iteratively updates large language models by mixing seed training data with filtered new outputs generated by the model itself. AMIE Tu et al. (2024) utilizes a self-play simulated learning environment for iterative improvement and mixes generated dialogues with supervised fine-tuning data through inner and outer self-play loops. SOTOPIA-$\pi$ Wang et al. (2024e) leverages behavior cloning from the expert model and self-generated social interaction trajectory to reinforce positive behaviors.
Another is generative replay, which adopts the self-generated synthesized data as knowledge to mitigate catastrophic forgetting. For instance, Self-Synthesized Rehearsal (SSR) Huang et al. (2024a) generates synthetic training instances for rehearsal, enabling the model to preserve its ability without relying on real data from previous training stages. Self-Distillation Fine-Tuning (SDFT) Yang et al. (2024b) generates a distilled dataset from the model itself to bridge the distribution gap between task datasets and the LLM’s original distribution to mitigate catastrophic forgetting.
@@ 6.1.2 Regularization-based
Regularization-based methods constrain the model’s updates to prevent significant deviations from original behaviors, exemplified by function- and weight-based regularization. Function-based regularization focuses on modifying the loss function that a model optimizes during training Zhong et al. (2023); Peng et al. (2023). For example, InstuctGPT Ouyang et al. (2022) employs a per-token KL-divergence penalty from the output probabilities of the initial policy model $\pi_{\text{SFT}}$ on the updated policy model $\pi_{\text{RL}}$. FuseLLM Wan et al. (2024) employs a technique akin to knowledge distillation Hinton et al. (2015), leveraging the generated probability distributions from source LLMs to transfer the collective knowledge to the target LLM.
Weight-based regularization Kirkpatrick et al. (2017) directly targets the model’s weights during training. Techniques such as Elastic Reset Noukhovitch et al. (2024) counters alignment drift in RLHF by periodically resetting the online model to an exponentially moving average of its previous states. Furthermore, Ramé et al. (2024) introduced WARM, which combines multiple reward models through weight averaging to address reward hacking and misalignment. Moreover, AMA Lin et al. (2024) adaptively average model weights to optimize the trade-off between reward maximization and forgetting mitigation.
@@ 6.1.3 Architecture-based
Architecture-based methods explicitly utilize extra parameters or models for updating, including decomposition- and merging-based approaches. Decomposition-based methods separate large neural network parameters into general and task-specific components and only update the task-specific parameters to mitigate forgetting. LoRA Hu et al. (2021); Dettmers et al. (2024) inject trainable low-rank matrices to significantly reduce the number of trainable parameters while maintaining or improving model performance across various tasks. This paradigm is later adopted by GPT4tools Yang et al. (2024a), OpenAGI Ge et al. (2024) and Dromedary Sun et al. (2024). Dynamic ConPET Song et al. (2023) combines pre-selection and prediction with task-specific LoRA modules to prevent forgetting, ensuring scalable and effective adaptation of LLMs to new tasks.
Merging-based methods, on the other hand, involve combining multiple models or layers to achieve general improvements, including but not limited to merging multiple generic and specialized model weights into a single model Wortsman et al. (2022); Ilharco et al. (2022); Yu et al. (2023a); Yadav et al. (2024), through mixture-of-expert approach Ding et al. (2024) or even layer-wise merging and up-scaling such as EvoLLM Akiba et al. (2024).
@@ 6.2 In-Context
MoT Li and Qiu (2023) Experience Insert
TRAN Yang et al. (2023b) Rationale Insert, Reflect
MemoryBank Zhong et al. (2024b) Experience, Rationale Insert, Reflect, Forget
MemGPT Packer et al. (2023) Experience Insert, Forget
TiM Liu et al. (2023a) Rationale Insert
IML Wang et al. (2024a) Rationale Insert, Reflect
ICE Qian et al. (2024) Rationale Insert, Reflect
AesopAgent Wang et al. (2024d) Experience, Rationale Insert, Reflect

### v2

@@ 4.2.2 Negative
In addition to acquiring positive solutions, recent research illustrates that LLMs can benefit from negative ones for self-improvement [140]. This strategy is analogous to trial and error in human behavior when learning skills. This section summarises typical methods of gaining negative solutions to assist in self-evolution.
@@
For instance, Self-Reward [148] generates responses, predicts corresponding rewards by the model itself, and constructs preference pairs based on rewards. SPIN [24] uses the results generated by the current model as negatives and the ground truth of SFT as positives for iterative self-training. Similarly, GRATH [21] generates both correct and incorrect answers. It then trains the model by comparing these two answers. Self-Contrast [153] contrasts the differences and summarizes these discrepancies into a checklist that could be used to re-examine and eliminate discrepancies. In ETO [105], the model interacts with the embodied environment to complete tasks and optimizes from the failure solutions. A3[141] improves ETO by adding rationale after each action for solving tasks. STE [120] implements trial and error where the model solves the tasks with unfamiliar tools. It learns by analyzing failed attempts to improve problem-solving strategies in future tasks. More recently, COTERRORSET [114] obtains incorrect solutions generated by PALM-2 and proposes mistake tuning, which requires the model to void making mistakes.
@@
Some methods add perturbation to generate harmful solutions [138, 70]. Given a task, RLCD [138] curates both positive and negative instructions and generates positive and negative solutions. DLMA [70] gathers both positive and negative instructional prompts and subsequently produces corresponding positive and negative solutions.
Rather than harmful perturbation, incorporating negative context is another way. Ditto [77] adds negative persona characters to generate incorrect conversations. The model then learns from the negative conversations to evolve persona dialogue ability.
@@ 4.3 Feedback
${\mathcal{F}}^{t}=f^{{\mathcal{F}}}({\mathcal{T}}^{t},{\mathcal{Y}}^{t},{\mathcal{E}}^{t},\mathrm{M}^{t};\mathrm{ENV}),$ (4)
@@ 4.3.1 Model
Current studies demonstrate that LLMs can play well as a critic [157]. In the cycle of self-evolution, the model judges itself to acquire the feedback of the solutions.
One type of feedback is a score that indicates correctness. Self-Reward [148], LSX Stammer et al [106], and DLMA [70] rate their own solutions and output the scores via LLM-as-a-Judge prompting. Similar to that, SIRLC [85] utilizes self-evaluation results of LLM as the reward for further reinforcement learning. Self-Alignment [155] leverages the self-evaluation capability of an LLM to generate confidence scores on the factual accuracy of its outputs.
Another type provides a textual description, offering multi-dimensional information. To alter the distribution of the responses via supervised learning, CAI [9] asks the model to critique its response according to a principle in the constitution. In contrast to supervised learning and reinforcement learning approaches, Self-Refine [80] allows the model to generate natural language feedback on its own output in a few-shot manner.
@@ 4.3.2 Environment
Another form of feedback comes from the environment, common in tasks where solutions can be directly evaluated. This feedback is precise and elaborate and can provide sufficient information for model updating. They may be derived from code interpreter [56, 22, 98], tool execution [90, 41], the embodied environment [13, 135, 164], and other LLMs or agents [125, 112, 118].
For code generation, Self-Debugging Chen et al [22] utilizes execution results on test cases as part of feedback while SelfEvolve [56] receives the error message from the interpreter. Similarly, Reflexion [98] also obtains the run-time feedback from the code interpreter. It then further reflects to generate thoughts. This run-time feedback contains the trace-back information that can point out the key information for improved code generation.
Recently, methods endow tool-using ability to LLMs and agents. Executing tools leading to feedback in return [41, 90, 105, 141, 120].
RoboCat [13] and SinViG [135] act in the robotic embodied environment. This type of feedback is precise and strong to guide self-evolution.
Communication feedback is common and effective in LLM-based multi-agent systems. Agents can correct and support each other, enabling co-evluation [125, 112, 118].
@@ 5 Experience Refinement
Figure 7: Experience refinement. $\tilde{{\mathcal{T}}}^{t}$ and $\tilde{{\mathcal{Y}}}^{t}$ are the task and solution after refinement, respectively. Filtering reduces the number of experiences from $n$ to $m$.
@@ 5.1.1 Metric-Based
By relying on feedback and pre-defined criteria, metric-based filtering improves the quality of the outputs [103, 90, 118, 127], ensuring the progressive enhancement of LLM capabilities through each iteration of refinement.
For example, ReSTEM [103] incorporates a reward function to filter the dataset sampled from the current policy. The function provides binary rewards based on the correctness of the generated samples rather than a learned reward model trained on human preferences in ReST [42]. AutoAct [90] leverages F1-score and accuracy as rewards for synthetic trajectories and collects trajectories with exactly correct answers for further training. Self-Talk [118] measures the number of completed subgoals to filter the generated dialogues, ensuring that only high-quality data is used for training. To encourage diversity of the source instructions, Self-Instruct [127] automatically filters low-quality or repeated instructions using ROUGE-L similarity and heuristics before adding them to the task pool.
@@ 5.1.2 Metric-Free
Some methods seek filtering strategies beyond external metrics, making the process more flexible and adaptable. Metric-Free filtering typically involves sampling outputs and evaluating them based on internal consistency measures or other model-inherent criteria [52, 129, 18]. The filtering in Self-Consistency [126] is based on the consistency of the final answer across multiple generated reasoning paths, with higher agreement indicating higher reliability. LMSI [52] utilizes CoT prompting plus self-consistency for generating high-confidence self-training data.
Designing internal consistency measures that accurately reflect output quality can be challenging. Self-Verification [129] allows the model to select the candidate answer with the highest interpretable verification score, calculated by assessing the consistency between the predicted and original condition values. For the code generation task, CodeT [18] considers both the consistency of the outputs against the generated test cases and the agreement of the outputs with other code samples.
@@ 5.2 Correcting
Recent advancements emphasize the importance of iterative self-correction in LLMs, allowing for the refinement of experiences. This section divides the methods employed into two categories: Critique-Based and Critique-Free correction. Critiques often serve as strong hints that include the rationale behind perceived errors or suboptimal outputs, guiding the model towards improved iterations.
@@ 5.2.1 Critique-Based
These methods rely on additional judging processes to draw the critiques of the experiences. Then, the experiences are refined based on the critiques. By leveraging either self-generated [80, 9, 98, 76] or environment-interaction generated critiques [41, 56, 164], the model benefits from detailed feedback for nuanced correction.
LLMs have demonstrated their ability to identify errors in their outputs. Self-Refine [80] introduces an iterative process in which the model refines its initial outputs conditioned on actionable self-feedback without additional training. To evolve from the correction, CAI [9] generates critiques and revisions of its outputs in the supervised learning phase, significantly improving the initial model. Applied to an agent automating computer tasks, RCI [57] improves its previous outputs based on the critique finding errors in the outputs.
Since weaker models may struggle significantly with the self-critique process, several approaches enable models to correct the outputs using critiques provided by external tools. CRITIC [41] allows LLMs to revise the output based on the critiques obtained during interaction with tools in general domains. SelfEvolve [56] prompts an LLM to refine the answer code based on the error information thrown by the interpreter. ISR-LLM [164] helps the LLM planner find a revised action plan by using a validator in an iterative self-refinement process.
@@ 5.2.2 Critique-Free
Contrary to critique-based, critique-free methods correct the experiences directly leveraging objective information [149, 22, 20, 40]. These methods offer the advantage of independence from nuanced feedback that critiques provide, allowing for corrections that adhere strictly to factual accuracy or specific guidelines without the potential bias introduced by critiques.
One group of critique-free methods modifies the experiences on the signal of whether the task was correctly resolved. Self-Taught Reasoner (STaR) [149] proposes a technique that iteratively generates rationales to answer questions. If the answers are incorrect, the model is prompted again with the correct answer to generate a more informed rationale. Self-Debug [22] enables the model to perform debugging steps by investigating execution results from unit tests and explaining the code on its own.
Different from depending on the task-solving signal, other information produced during the solving process can be leveraged. IterRefinement [20] relies on a series of refined prompts that encourage the model to reconsider and improve upon its previous outputs without any direct critique. For information extraction tasks, Clinical SV [40] grounds each element in evidence from the input and prunes inaccurate elements using supplied evidence.
@@ 6.1 In-Weight
Classical training paradigms in updating LLMs in weight encompass continuous pretraining [14, 93], supervised fine-tuning [75], and preference alignment [83, 115]. However, in the iterative training process of self-evolving, the core challenge lies in achieving overall improvement and preventing catastrophic forgetting, which entails refining or acquiring new capabilities while preserving original skills. Solutions to this challenge can be categorized into three main strategies: replay-based, regularization-based, and architecture-based methods.
@@ 6.1.1 Replay-based
Replay-based methods reintroduce previous data to retain old knowledge. One is experience replay, which mixes the original and new training data to update LLMs [93, 141, 156, 60, 124]. For example, Reinforced Self-Training (ReST) [42, 4] method iteratively updates large language models by mixing seed training data with filtered new outputs generated by the model itself. AMIE [117] utilizes a self-play simulated learning environment for iterative improvement and mixes generated dialogues with supervised fine-tuning data through inner and outer self-play loops. SOTOPIA-$\pi$ [125] leverages behavior cloning from the expert model and self-generated social interaction trajectory to reinforce positive behaviors.
Another is generative replay, which adopts the self-generated synthesized data as knowledge to mitigate catastrophic forgetting. For instance, Self-Synthesized Rehearsal (SSR) [53] generates synthetic training instances for rehearsal, enabling the model to preserve its ability without relying on real data from previous training stages. Self-Distillation Fine-Tuning (SDFT) [143] generates a distilled dataset from the model itself to bridge the distribution gap between task datasets and the LLM’s original distribution to mitigate catastrophic forgetting.
@@ 6.1.2 Regularization-based
Regularization-based methods constrain the model’s updates to prevent significant deviations from original behaviors, exemplified by function- and weight-based regularization. Function-based regularization focuses on modifying the loss function that a model optimizes during training [160, 87]. For example, InstuctGPT [83] employs a per-token KL-divergence penalty from the output probabilities of the initial policy model $\pi_{\text{SFT}}$ on the updated policy model $\pi_{\text{RL}}$. FuseLLM [119] employs a technique akin to knowledge distillation [46], leveraging the generated probability distributions from source LLMs to transfer the collective knowledge to the target LLM.
Weight-based regularization [58] directly targets the model’s weights during training. Techniques such as Elastic Reset [82] counters alignment drift in RLHF by periodically resetting the online model to an exponentially moving average of its previous states. Furthermore, Ramé et al [92] introduced WARM, which combines multiple reward models through weight averaging to address reward hacking and misalignment. Moreover, AMA [69] adaptively average model weights to optimize the trade-off between reward maximization and forgetting mitigation.
@@ 6.1.3 Architecture-based
Architecture-based methods explicitly utilize extra parameters or models for updating, including decomposition- and merging-based approaches. Decomposition-based methods separate large neural network parameters into general and task-specific components and only update the task-specific parameters to mitigate forgetting. LoRA [50, 29] inject trainable low-rank matrices to significantly reduce the number of trainable parameters while maintaining or improving model performance across various tasks. This paradigm is later adopted by GPT4tools [139], OpenAGI [39] and Dromedary [108]. Dynamic ConPET [104] combines pre-selection and prediction with task-specific LoRA modules to prevent forgetting, ensuring scalable and effective adaptation of LLMs to new tasks.
Merging-based methods, on the other hand, involve combining multiple models or layers to achieve general improvements, including but not limited to merging multiple generic and specialized model weights into a single model [130, 55, 147, 137], through mixture-of-expert approach [33] or even layer-wise merging and up-scaling such as EvoLLM [3].
@@ 6.2 In-Context
MoT [66] Experience Insert
TRAN [140] Rationale Insert, Reflect
MemoryBank [161] Experience, Rationale Insert, Reflect, Forget
MemGPT [84] Experience Insert, Forget
TiM [72] Rationale Insert
IML [121] Rationale Insert, Reflect
ICE [89] Rationale Insert, Reflect
AesopAgent [123] Experience, Rationale Insert, Reflect